OpenStudy (anonymous):

find \(cos^{-1}(2)\)

OpenStudy (anonymous):

make graph of arccos(x)

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=arccos%282%29

OpenStudy (anonymous):

its in comple , cant graph it

OpenStudy (anonymous):

i need to how convert it :-\ dnt need only answer :(

OpenStudy (anonymous):

that is like cos x =2 how would cos x exist when it lies only in range [-1,1] ??????

OpenStudy (anonymous):

complex *

ganeshie8 (ganeshie8):

cosy = 2

OpenStudy (anonymous):

Oh ok

ganeshie8 (ganeshie8):

e^iy + e^(-iy) = 4

OpenStudy (anonymous):

ok then ?

ganeshie8 (ganeshie8):

no clue, im here to learn how to solve this from you :)

OpenStudy (anonymous):

Is the answer 1.32i

OpenStudy (anonymous):

ok i might continue mmm ************* the funny thing is a car was going to hit me , but God it stoped at the last sec ! when the driver stoped he looked at me and said are u oki miss :O i said no :( idk how to solve arccos 2 :'( haha :P ********************* ok ill try

OpenStudy (anonymous):

ok i got it \(cos ^{-1} z= -i[z+i(1-z^2)^\frac{1}{2}\)] mmm

OpenStudy (anonymous):

Do you know \[\huge \cos(z)=\frac{ e ^{iz} + e ^{-iz} }{ 2 }\]

OpenStudy (anonymous):

ik it mmm but im trying to derive arccos from it

OpenStudy (anonymous):

Multiply by \[\huge e ^{iz}\]

OpenStudy (anonymous):

but since i got the formula :D thx all :D

OpenStudy (anonymous):

And solve the quadratic

hartnn (hartnn):

e^iy + e^(-iy) = 4 x+1/x = 4 x = \(2 \pm \sqrt3\) e^iy = x y = - i log x

OpenStudy (anonymous):

and take the log u will get the answer as 1.32i

hartnn (hartnn):

y = - i ln x

OpenStudy (anonymous):

sry i made a typo \(cos ^{-1} z= -i\log [z+i(1-z^2)^\frac{1}{2}]\) \(cos ^{-1} z= -i\log [2+i(1-2^2)^\frac{1}{2}]\) \(cos ^{-1} z= -i\log [ 2-\sqrt 3]\)

OpenStudy (anonymous):

the hartnn answer :D ty !